Harry can paint a room in 3 hours, and kerry can paint it in 4 hours. how long will it take if they work together?

Answer by unlockmath(1599)

Hello,
Let x be the number of hours they can both do the work. The equation weill be set up as follows:
1/3 + 1/4 = 1/x Multiply by 12x to give us:
4x+3x=12 Combine and divide each side by 7 to get:
x=12/7 hours or rewritten as 1 5/7 hours
RJ
There is a book that might help:
www.math-unlock.com